# Following Is All You Need: Robot Crowd Navigation Using People As Planners

### v1

@@ IV-B Quantitative Evaluation
In comparison, ORCA tends to keep a close distance from humans since it assumes reciprocal avoidance behaviours, which lead to more collisions than ours. SF and DWA are less robust in highly dynamic environments, as shown by the inconsistent performance across the three scenes. For Pred2Nav, although SGAN gives more accurate human predictions on large-scale datasets [43], it doesn’t consistently improve the planning performance. HEIGHT is less sensitive to collisions because it is not explicitly enforced like the other methods. It demonstrates biased behaviours such as moving in a curved path, which results in heavy collisions with the wall in the Promenade scene. We also show that by simplifying complex crowd navigation problems using people as planners, we can achieve the best performance using a simple base planner SF, which is less capable when planning independently.
@@ IV-C Qualitative Evaluation
Using people as planners, our method demonstrates efficient and socially-aware behaviours that contribute to the outstanding results presented above. For the Promenade scene in Fig. 5-1a, the robot switched to the second leader when the first leader was temporarily occluded, and switched back to the first leader when the second leader stopped its movement. In Fig. 5-1b, Pred2Nav resulted in collisions twice. From the sharp turns before the collisions, we can infer that the robot tried to avoid the humans, but it acted too late. Compared to Fig. 5-1c, predicting with SGAN did not bring much benefit. For the Crossing scene, both SF and HEIGHT resulted in collisions with a fast-moving bicycle, as shown in Fig. 5-2b and Fig. 5-2c. Our method, as shown in Fig. 5-2a, avoided this encounter by switching between the two leaders.
For the Roundabout scene, DWA first collided with a bicycle, then with two humans walking from the opposite direction, as shown in Fig. 5-3b. In Fig. 5-3c, ORCA only collided with the bicycle, but was forced to stop by the two humans and only continued moving after the humans walked by. In comparison, as shown in Fig. 5-3a, our method followed the leader to cross the road safely, switched to default planner in the middle when the leader was temporarily occluded, and then continued to follow towards the goal. An interesting observation in this scene is that, although traffic rules are not explicitly defined, our method follows the human to cross the road from the appropriate point on the sidewalk, which avoids a dangerous encounter with the bicycle. Other baselines, however, move towards the goal directly, which inevitably increases the difficulty of avoiding fast-moving road users on the main road. This makes our method especially useful in real-world deployments, as it is difficult to define every rule in every scenario, our method can simply follow the surrounding agents to respect the “unknown rules”.
@@ IV-D Real-World Experiment
[…]
@@ V CONCLUSIONS
This paper presents a novel crowd navigation scheme which utilises people as external intelligent planners. The more complex planning decisions are handed over to humans so the robot’s task is simplified into straightforward subgoal planning. The experiments demonstrate robot behaviours that are safe, efficient, and socially compliant in various crowded environments. Our proposed framework offers new insights into human-robot relationships in social navigation. In a rare scenario when a dense crowd walk against the robot, our method will be reduced to the base planner as no suitable leader can be identified. It then relies on the base planner’s capability to avoid or retreat. Future work can be done to infer useful information from other humans who are not leading the robot.

### v2

@@ IV-B Quantitative Evaluation
In comparison, ORCA tends to get too close to humans since it assumes reciprocal avoidance behaviours, which lead to more collisions than ours. SF and DWA are less robust in highly dynamic environments, as shown by the inconsistent performance across the three scenes. For Pred2Nav, although SGAN gives more accurate human predictions on large-scale datasets [45], it doesn’t consistently improve the planning performance. HEIGHT is less sensitive to collisions because it is not explicitly enforced like the other methods. It demonstrates biased behaviours such as moving in a curved path, which results in heavy collisions with the wall in the Promenade scene. We also show that by simplifying complex crowd navigation problems using people as planners, we can achieve the best performance using a simple base planner SF, which is less capable when planning independently.
@@ IV-C Qualitative Evaluation
Using people as planners, our method demonstrates efficient and socially-aware behaviours that contribute to the outstanding results presented above. For the Promenade scene in Fig. 5-1a, the robot switched to the second leader when the first leader was temporarily occluded, and switched back when the second leader stopped moving. In Fig. 5-1b, Pred2Nav resulted in collisions twice. From the sharp turns before the collisions, we can infer that the robot tried to avoid the humans, but it acted too late. Compared to Fig. 5-1c, predicting with SGAN did not bring much benefit. For the Crossing scene, both SF and HEIGHT resulted in collisions with a fast-moving bicycle, as shown in Fig. 5-2b and Fig. 5-2c. Our method, as shown in Fig. 5-2a, avoided this encounter by switching between the two leaders.
For the Roundabout scene, DWA first collided with a bicycle, then with two humans walking face-to-face, as shown in Fig. 5-3b. In Fig. 5-3c, ORCA only collided with the bicycle, but was forced to stop by the two humans and only continued moving after the humans walked by. In comparison, as shown in Fig. 5-3a, our method followed the leader to cross the road safely, switched to default planner in the middle when the leader was temporarily occluded, and then continued to follow towards the goal. An interesting observation in this scene is that, although traffic rules are not explicitly defined, our method follows the human to cross the road from the appropriate point on the sidewalk, which avoids a dangerous encounter with the bicycle. Other baselines, however, move towards the goal directly, which inevitably increases the difficulty of avoiding fast-moving road users on the main road. This makes our method especially useful in real-world deployments, as it is difficult to define every rule in every scenario, our method can simply follow the surrounding agents to respect the “unknown rules”.
@@ IV-D Real-World Experiment
In addition to the performance comparison against baselines using recorded human data, we further demonstrate the effectiveness of our method when interacting with real-world reactive humans. We conduct the experiment in a 150m crowded campus corridor with over 40 human encounters, which is similar to the previous SDD scenes in terms of environment and human density. We use a 2D LiDAR to obtain point clouds for reachability check, and use a depth camera for human tracking. To ensure human safety and privacy, we reduce the robot speed limit to $1.0\text{m/s}$, stop the robot when getting too close to pedestrians, and do not collect their identifiable information during experiments.
@@
Figure 7: Failure cases in simulation experiments. In (1), the robot was confused by a human blocking the way and forced into a short deadlock. In (2a), the robot was following a group and did not detect an oncoming bicycle due to occlusion. A dataset labelling error caused the bicycle to unrealistically run over the pedestrians, leading to a sudden collision with the robot in (2b).
IV-E Limitations
During simulation experiments, we found that our method may fail when interacting with aggressive road users, as shown in Fig. 7. Future work could incorporate data-driven policies trained on real-world corner cases. Some existing works allow the robot to actively influence human behaviours [46, 47, 48], which may also be useful, especially in difficult scenarios such as moving against a dense crowd.
The dynamic subgoal setting in the proposed framework may potentially affect trajectory smoothness. In the current experiments, our method does not show significant fluctuations compared to the baselines (comparison results on trajectory smoothness can be found in the supplementary video). Nevertheless, to improve the smoothness, future work may consider other types of robot dynamics apart from the holonomic model, or incorporate various trajectory smoothing techniques [49].
@@ V CONCLUSIONS
This paper presents a novel crowd navigation scheme which utilises people as external intelligent planners. The more complex planning decisions are handed over to humans so the robot’s task is simplified into straightforward subgoal planning. The experiments demonstrate robot behaviours that are safe, efficient, and socially compliant in various crowded environments. Our proposed framework offers new insights into human-robot relationships in social navigation.